### v1

# On deformation quantization of the space of connections on a two manifold and Chern Simons Gauge Theory

Abstract.
We use recent progress on Chern-Simons gauge theory in three dimensions [18] to give explicit, closed form formulas for the star product on some functions on the affine space of (smooth) connections on the trivialized principal -bundle on a compact, oriented two manifold These formulas give a close relation between knot invariants, such as the Kauffman bracket polynomial, and the Jones and HOMFLY polynomials, arising in Chern Simons gauge theory, and deformation quantization of This relation echoes the relation between the manifold invariants of Witten [20] and Reshetikhin-Turaev [16] and geometric quantization of this space (or its symplectic quotient by the action of the gauge group). In our case this relation arises from explicit algebraic formulas arising from the (mathematically well-defined) functional integrals of [18].
***PRELIMINARY***
Contents
1. Introduction
In [2] Bayen, Flato, Fronsdal, Lichnerowicz, and Sternheimer proposed deformation quantization as a method of constructing algebraically the effect of quantization by deforming the algebra of functions on a symplectic manifold or Poisson manifold They conjectured, and in some cases proved, that on such a manifold there exists an associative product on the formal power series with the property that
for where is the Poisson bracket on
The construction of the star product for symplectic manifolds in general was given by deWilde and Lecomte [6] and by Fedosov [7], and for general Poisson manifolds by Kontsevich [15]. The resulting star products are often not so easy to compute.
The purpose of this paper is to use recent progress on Chern-Simons gauge theory in three dimensions [18] to give explicit, closed form formulas for the star product on some functions on the affine space of (smooth) connections on the trivialized principal -bundle on a compact, oriented two manifold These formulas give a close relation between knot invariants arising in Chern Simons gauge theory and deformation quantization of echoing the relation between the manifold invariants of Witten [20] and Reshetikhin-Turaev [16] and geometric quantization of this space (or its symplectic quotient by the action of the gauge group). In our case this relation arises from explicit algebraic formulas in the quantum field theory of [18], which is mathematically well defined.
Let be a compact, oriented two manifold, and a compact Lie group.111We will later generalize this to the case of noncompact classical groups. Choose a representation of so that the resulting trace gives a symplectic form on the space of connections on the trivial principal -bundle on In more detail, if we use the trivialization to identify the symplectic form is given in [1] by
where and The explicit formula shows that the symplectic form is invariant under affine translations on 222The symplectic form is also invariant under the action of the gauge group. This action does not commute with the affine translations, and this fact will have consequences we will study later in the paper. See Section 5.2.
The Poisson bivector field on can be written down in local coordinates; we have
where we choose a basis for and write
Since is an affine space, and since is translation invariant, the star product can be computed using formulas discovered by [2] in their original paper333See also [7], p. 222. Explicitly, denote by the Poisson bivectorfield associated to the symplectic form Then
where denotes the multiplication map on functions.
The expression (1.2) has the appearance of a canonical object, but in fact it involves an implicit choice. The reason is that Fedosov’s construction of the star product depends on a choice of a symplectic connection on the underlying manifold; that is, a torsion-free connection on the tangent bundle preserving the symplectic form. In the case of an affine space, there is a natural choice of such a connection, and it leads to the formula (1.2). But there are other choices; in particular the choice of the affine connection on is not a gauge invariant choice. We will see that this has consequences for the behavior of the star product under reduction by the action of the gauge group.
The affine space is equipped with some functions (the Wilson loops) arising from curves on Choose a representation of (which may or may not be the same as the one used to define the symplectic form). Given a closed, oriented curve in define by
where is any basepoint on and is the trace in the representation (To simplify the notation, we will suppress the basepoint and the representation when possible.)
The main result of this paper will be a computation of the star product on functions generated by the we have to impose intersection conditions on the curves appearing in any given star product. In view of (1.2), the first task is to compute the Poisson bracket
of two such functions, which from now on we assume to intersect transversally. Consider the case where the representations for both the symplectic form and the Wilson loops are given by the defining representation, and two curves intersect at one point We will show that
where is the curve obtained by concatenating the oriented curves and at their intersection point 444The star notation is used in the literature to denote both the star product in deformation quantization, and the concatenation of loops. We have maintained this usage. I do apologize. More generally, we will show (see (2.12)) that if there is more than one intersection point, the Poisson bracket is given by the sum of the Wilson loops obtained from concatenating the two curves at each of the intersection points. We will obtain similar formulas for the Poisson bracket for the groups and though these involve unoriented curves.
In all these cases our formulas for the Poisson bracket coincide with those discovered by Goldman [10] for the Poisson bracket of the Wilson loops, considered as functions on the Marsden-Weinstein quotient of the space of connections by the action of the gauge group This space is the character variety of conjugacy classes of representations of the fundamental group of in The reason for the agreement between our formulas and Goldman’s is that the both the functions and the symplectic form, and hence the Poisson bracket, are gauge invariant; the identity between Poisson brackets of invariant functions on a Hamiltonian -space and the corresponding Poisson brackets on the reduced space follows by general principles of symplectic geometry. This is not the case for star products, unless additional conditions are imposed.
In addition to the resemblance to Goldman’s results, there is also a close similarity between the Poisson bracket (1.3) and formulas arising in Chern-Simons gauge theory [18], generalized from the case of to that of Consider the space of connections on the trivialized prinicipal -bundle on If we impose the axial gauge condition the Chern-Simons action on the resulting space of compactly supported connections is given by
The propagator for this theory is obtained by inverting the differential operator in the quadratic action and is given by [18]
Here and we have chosen a basis for and define 555The trace is taken in the representation used to define the Chern Simons function. likewise we choose an oriented basis for the fibres of the cotangent bundle at a point and its pullback to agreeing with the orientation of Finally666The addition of a constant to although in principle allowable, results in no change to any expectation, due to the fact that is skew.
We note that there is a resemblance between the propagator (1.4) and the differential operator appearing in the Poisson bivector field (1.3). To exploit the similarity, consider the operator
or, more explicitly
The operator is the operator which evaluates expectations in Chern-Simons gauge theory. And we have, and for any two curves in for any and any
Theorem 1.
For any two curves in intersecting transversally, for any and any
with similar formulas for the Poisson brackets of functions given by polynomials in Wilson loops.
Here for any curve in the curve denotes the curve in obtained by translating by in the vertical direction777We consider self crossings of the individual curves and as virtual crossings, to which the operators are not to be applied. , and is the projection. Likewise
Theorem 2.
For any two curves in intersecting transversally, for any and any
with similar formulas for the Poisson brackets of functions given by polynomials in Wilson loops.
Thus we may define the star product on using expectations in Chern-Simons gauge theory, that is
Definition 1.9.
For any two curves in for any and any define
In Section 4 we will prove888Since is an infinite dimensional space, and since our star product is not defined everywhere, this does not quite follow from the results of [2, 7].
Theorem 3.
The star product is associative where defined; that is, if are curves in which intersect pairwise transversally, then
This relation between deformation quantization of the space of connections on a two manifold and Chern-Simons gauge theory on is the main result of this paper. Now in [18] we found explicit formulas for evaluating the expectations of Wilson loops in Chern-Simons gauge theory in axial gauge, such as the expressions
The resulting formulas turned out to be close relatives of the Jones and HOMFLY polynomials. We therefore obtain as a consequence of this relation explicit formulas, related to the Jones and HOMFLY polynomials, for the star product of Wilson loops on the space of connections (See Section 4.1). These arise from the formulas derived in [18] for expectations in Chern Simons gauge theory; explicitly, these are formulas (3.7) and (3.8) (in the case of and ) or formulas (3.13) and (3.14) (in the case of or )
In view of the fact that the Poisson bracket of Wilson loops on the space of connections agreed with Goldman’s formula for the Poisson bracket of those functions, considered as functions on the moduli space of flat connections, it is natural to ask whether a similar phenomenon occurs for the star product, and whether our formulas give a deformation quantization of these moduli spaces; in the case of we would then obtain a deformation quantization of Teichmuller space. And in [8] Fedosov shows that, in the case of finite dimensional symplectic manifolds, deformation quantization commutes with reduction, and that in an appropriate sense, the star product on a Hamiltonian -space, restricted to invariant functions, gives a star product on the symplectic quotient.
However a closer look at the formulas formulas (3.7) and (3.8) (in the case of and ) and (3.13) and (3.14) (in the case of or ) shows that the star product of two Wilson loops is not invariant under homotopy. This causes no difficulty for the star product on , but on the symplectic quotient, homotopic loops give rise to the same function on the moduli space of flat connections, so that the star product must be homotopy invariant to be well defined.
The failure of homotopy invariance was already noted in [18] as the failure of the expectations in Chern-Simons gauge theory to be invariant under the second Reidemeister move. In [18] we showed that, for this could be remedied by using a background gauge field to shift the value of the expectation of a single unknotted circle from its expected value in in gauge theory to the value expected999Here is a parameter appearing both in Kauffman’s bracket polynomial and in Chern-Simons gauge theory. See [13] and [18]. from Kauffman’s bracket polynomial [13].101010In [18] there are normalization conventions slightly different from Kauffman’s which adjust these expectations by a sign. From the point of view of Fedosov’s construction, the fact that the star product on does not descend in a direct way to the quotient may be expected due to the fact that our construction of the star product used implicitly the affine symplectic connection on which is not gauge invariant. We would hope that this point of view may be helpful also for Chern-Simons gauge theory. In particular it would be interesting to see whether a more sophisticated handling of the time independent gauge transformations in Chern-Simons gauge theory in axial gauge may give insight into the rather mysterious shift
The structure of this paper is as follows. In Section 2 we repeat the computation from [18] of the Poisson bracket of functions given by polynomials in Wilson loops, giving formulas that agree with those computed by Goldman on the moduli space of flat connections. In Section 3 then recall the results of [18], and extend them to the manifold this extension is straightforward. In Section 4 we compare Chern-Simons functional integral on with deformation quantization of and show that the star product of Wilson loops arising from transversally intersecting curves, considered as functions on may be computed using the expectations defined and computed in [18]. We state this as Definition 4.4. We then prove that associativity of the star product, where defined, follows from properties of expectations in the linear functional corresponding to Chern-Simons functional integral. We close the section with an algorithm which gives an explicit formula for the star product of Wilson loops. Finally in Section 5, we briefly review the deformation quantization of Kahler manifolds arising from geometric quantization in the works of Boutet de Monvel and Guillemin [4], Bordemann, Meinrenken and Schliechenmeier [3] and Guillemin [11]. We show how applying these results to the moduli space of flat connections on a Riemann surface hints at a relation between deformation quantization of the moduli space of flat connections on a Riemann surface and the topological quantum field theory arising from geometric quantization of that space. This parallels the relation we found in this paper between deformation quantization of the space of connections and Chern-Simons gauge theory. We close by discussing the relation between the (non)-invariance of the expectations in the Chern-Simons gauge theory of [18] under the Second Reidemeister move and the (non)-invariance of the star product on under reduction. While the lack of invariance of expectations in Chern-Simons gauge theory is mysterious, the non-invariance of the star product arises from a known cause—the gauge non-invariance of the affine symplectic connection on So this hints at a possible resolution of both issues, as mentioned above.
2. The Poisson Bracket
As a first step toward computing the star product on we compute the Poisson bracket of two Wilson loops on this space. We perform these computations by methods similar to those used in [18] for expectations in Chern-Simons gauge theory; in Section 4 we will see that this apparent analogy is in fact an identity. We will see that these computations give results identical to those given by Goldman [10] on the moduli space of flat connections on the two manifold. Since this moduli space is the reduced space of the space of connections by the action of the gauge group, and since the Wilson loops are invariant under this group, this is to be expected.
Let be a closed, oriented two manifold, let be a Lie group as before, along with a choice of representation whose associated trace gives a nondegenerate inner product on and let be the space of connections on the trivialized principal -bundle on Then This space is equipped [1] with a symplectic form given by
where and and the trace is the trace in the chosen representation.
The Poisson bivector field arising from the symplectic form is given in local coordinates by
Here we have chosen a basis for and write
Remark 2.2.
An invariant expression for is given as follows. Given a differentiable function on the derivative is The orientation of gives a map given by for In these terms, morally for any differentiable functions on whose derivatives are in the image of (In general, these derivatives may correspond under to distributional one forms.) This coordinate free definition agrees with (2.1).
Suppose that we are given two oriented simple closed curves in intersecting transversally at a finite number of points We wish to compute the Poisson bracket
(where as before we write for the trace in some representation of the holonomy of the connection along the oriented path given by with basepoint and suppress for notational simplicity the basepoint where we are taking a trace, as well as the representation) by applying the skew operator to the pair of functions
We obtain
where the are the transverse intersection points of and
To perform computations of this type, we recall some formulas from [18].
We begin by studying connections on the circle. Let be a connection on the trivial principal bundle on with its standard orientation, and let We compute
Here we have written for the holonomy of the connection along the circle from the point to itself in the direction given by the standard orientation, and we consider both and as subsets of some space of matrices corresponding to the chosen representation.
In terms of functional derivatives, this equation (2.4) can be written
where is the projection arising from the representation we chose for (inside some ) and the inner product arising from the corresponding trace.
We also record this functional derivative for the case of a connection on We have (using again the representation in )111111The factor of in the equations below is due to the derivative being taken at the endpoint, and can be seen by applying the analog of (2.4) to a series of approximations of the delta function by smooth functions. If the connection on a circle is thought of as being a connection on the interval with endpoints identified, the factors of arising from the endpoints add to give (2.4).
Likewise
Given an embedded path, or curve, in a manifold, and a connection on the trivialized principal bundle on the manifold, these formulas give expressions for the derivatives in the tangent direction to the curve of the holonomy of the pulled back connection.
We now return to (2.3). Applying (2.5), the contribution of the -th point to this sum is121212Strictly speaking, we would need to smooth the functions To do this, smooth the connection by convolving with a delta sequence and write (where here denotes convolution). Applying to the product of two such functions and taking the limit we obtain (2.8).
where, as before, is a basis for and is a sign arising from any symplectic form on giving the orientation of evaluated on the tangent vectors to the oriented curves; explicitly, if are tangents to the oriented curves at pointing in the direction of the orientations, then131313Note that if the curves do not intersect, the corresponding Wilson loops Poisson commute; see [19] for a proof of this fact along the lines of the gauge theoretic computation given here.
where is a symplectic form on the oriented two manifold with
If we choose the same representation for the Wilson loops as we did in defining the symplectic form, so that we have
where is the map given by composing the inclusion of into given by the chosen representation, followed by the restriction to of the projection in the metric arising from the trace
In the case where and the representation is the defining representation, the map is the identity, and we obtain
where denotes the concatenation of the oriented curves and at the common basepoint This agrees with Goldman’s computation (Theorem 3.13 of [10]).
In the case of and the defining representation, the map is given by
(cf. Corollary 1.10 of [10]). We therefore have
where we denote by the curve with the orientation reversed. This is in agreement with Theorem 3.17 of [10].
Note that the Poisson bracket we have computed on extends immediately (by the Leibniz rule) to functions given by polynomials in Wilson loops, so long as they obey (pairwise) the transverse intersection condition.
Remark 2.14.
It is not so clear that the computation of the Poisson bracket on (nicely intersecting) Wilson loops suffices to determine its value on all smooth functions on or on the moduli space of flat connections. In finite dimensions, Nachbin’s Theorem gives a criterion for an algebra of functions to be dense in the algebra of smooth functions on a smooth manifold:
Theorem 2.15 (Nachbin [14]).
Let be a subalgebra of the algebra of smooth functions on a finite dimensional smooth manifold Suppose that
For each there exists so that
For each pair with there exists so that
For each and each with there exists so that
Then is dense in the Frechet space
In the case of smooth moduli spaces, this theorem can be applied. But the algebra generated by Wilson loops given by the defining representation of does not separate tangent vectors at the equivalence class of the trivial representation,141414This representation occurs only in a singular quotient of so this argument is not quite on all fours, but is indicative of the issues to be dealt with. due to the fact that elements of are traceless in this representation. Similar problems occur for other groups. It is possible that including Wilson loops in all representations remedies this problem.
Remark 2.16 ( and ).
Note that as in the case of , the Poisson bracket of Wilson loops in the fundamental representation of or is obtained from formula (2.10) with
Thus in all three cases we obtain
and therefore
When (and also ), this expression coincides with (2.13), since for or
so that
See [10].
3. Chern Simons gauge theory in axial gauge on
Let be a compact, oriented two manifold, let be a Lie group, and consider the space of connections on the trivialized principal -bundle on Choose a representation of ; this gives a trace on the lie algebra The space of connections may be identified with and in these terms the Chern-Simons invariant of a connection is given by
As in [18], let denote the subspace of the space of connections satisfying the axial gauge condition
where is the coordinate on arising from projection to
The Chern-Simons invariant of an element is then151515We absorb a factor of into the coupling in the functional integrals.
As in [18], we consider the formal functional integral with action Expectations in this formal functional integral can be computed using the propagator, the formal inverse of the operator appearing in the definition of and given by (1.4), viz
Here we have chosen a basis for and define 161616The trace is taken in the representation used to define the Chern Simons function. Also and171717The addition of a constant to although in principle allowable, results in no change to any expectation, due to the fact that is skew.
The operator gives well defined formulas, as in [18], for the expectations of polynomials in the gauge field and these can be thought of as giving a linear functional on the algebra of such polynomials.181818Again, as in Remark 2.2, we may write the operator in a form independent of choice of coordinates on showing in particular that it depends only on an orientation on not on the choice of local coordinates. We wish to find a way to define expectations of polynomials in Wilson loops, which are not of this form. We proceed as in [18].
Let be closed, oriented, immersed curves in , whose projections to intersect pairwise transversally at a finite number of points of with inverse images As in [18], the formal functional integral
for should be taken to be
where the is a linear functional on the vector space generated by monomials in the involving curves satisfying our intersection conditions and given by
where
Note that the linear functional which is defined on a space of functions on the space of connections, involves only functional differentiation, not functional integration.
Note also that the definition of does not contain terms referring to possible self-intersections of the curves for reasons that will become clear when we compute the effect of the Poisson bivector field. In the language of [13] and [18], we consider these self-intersections to be virtual intersections.
In [18] we developed formulas to compute the value of a linear functional of the type in the case where was replaced by These same formulas carry over immediately to the case we are considering here, and can be summarized, using the standard diagrams in knot theory, as follows.
Consider first the case where and where we use the defining representation for both the Chern-Simons functional and the Wilson loops. Then we showed
We then showed how to compute the exponentials of and proved
and
where
Here we have used the notation () for a ”bare” crossing; one where all the appropriate operators have already been applied. This notation has been used by Kauffman [13] to denote a virtual crossing. Similarly () for the ”honest” smoothed intersection (in our case involving a ”jump”).
This formula applies also to the groups and again in the defining representation.
In the case where or , the Wilson loop is independent of orientation, since for or
We can then express formulas (3.5) and (3.6) (after a normalization ) in terms of unoriented curves as
and
where
and
Likewise, if we take or we have the formulas
We may again exponentiate191919There are subtleties involved in computing the action of on the smoothings, which involve jumps, appearing in (3.11) and (3.12). See [18] for the detailed computations. to obtain
while
where
Note that if we set equations (3.13) and (3.14) differ from their counterparts (3.5) and (3.6) only by factors of corresponding to knot framing.
All these formulas are taken verbatim from [18]. In the next section, we will see how they give a star product on the space of connections on
4. Chern Simons gauge theory in three dimensions and the star product on the space of connections in two dimensions
In this Section we show that the uncanny resemblance between the formulas (2.12) and (2.13) obtained for the Poisson bracket and the formulas (3.11) - (3.12) and (3.3) - (3.4) for the the action of the operator in both the and cases are not coincidences; they arise because there is an identity between Poisson brackets of Wilson loops in two dimensional gauge theory and expectations of related objects in Chern-Simons gauge theory. This identity echoes the relation between the manifold invariants of Witten and Reshetikhin-Turaev and geometric quantization of moduli spaces.
To do this we compare the operators and
Recall that as in (1.6), the operator is given by
On the other hand, the bivector field is given (1.1) by the operator
Hence, if are two curves in intersecting transversally at a finite number of points in and if we denote by the translation by in the direction in we have, for any and any
where is the projection and denotes multiplication in the algebra of functions on We summarize these results below.
Theorem 1.
Let and be two polynomials in Wilson loops on corresponding to curves which intersect pairwise transversally. Then
Here the operator is defined on polynomials in the Wilson loops by defining
for each monomial in the Wilson loops, and extending the definition to polynomials by linearity.202020The factor of in (4.3) is due to the difference between the action (by the product rule) of the symmetric second order differential operator on the product of the two functions and the contraction of the skew bivector field with
It follows that
Theorem 2.
Let and be two polynomials in Wilson loops on corresponding to curves which intersect pairwise transversally. Then
where we consider all intersections of the loops appearing in to be virtual crossings, likewise all intersections of loops appearing in
Now, on an affine space like the formula of [2] for the star product (in terms of Fedosov’s construction, the star product associated with the symplectic connection invariant under affine translations; see [7], p. 222) is
Definition 4.4.
The star product of two polynomials in Wilson loops that intersect pairwise transversally is given by (setting )
We verify directly that this product is associative where defined:
Theorem 3.
The star product is associative where defined; that is, if and are polynomials in Wilson loops in corresponding to curves which intersect pairwise transversally, then
Proof.
We have (recalling we set )
while
But, since depends only on the sign of
4.1. Explicit formulas for the star product
Definition 4.4, combined with formulas (3.7) and (3.8) (in the case of and ) or formulas (3.13) and (3.14) (in the case of or ) give explicit formulas for the star product:
As a first example, consider the case where Let be two immersed curves in and suppose intersect transversally at one point and that the tangents to the oriented curves at satisfy for some symplectic form on with Then applying Theorem 2 and equation (3.5), and recalling we have
Note that for small we have, comparing with (2.17),
as expected.
More generally, the algorithm for computing the star product of any two monomials and can be given as follows.
Step 1. Translate by in the vertical direction to obtain a collection of curves in Likewise translate by to obtain another collection of curves in Consider the product
Step 2. For any pair , and any intersection point remove and from the product and replace it with
where
and where are given as in (3.7) and (3.8) (in the case of and ) or formulas (3.13) and (3.14) (in the case of or ).
In other words, replace by
Step 3. Repeat Step 2 for all other intersection points.212121The formula for the kind of Wilson loop arising from concatenation of two curves is not quite the same as that in (4.7) when one of the curves is itself a concatenation; it is the trace of a product of holonomies along segments. We omit the precise formulas since the geometric picture is so clear.
5. Some Comments and Speculations: Relation to Geometric Quantization; Quantization of Moduli spaces
5.1. Geometric Quantization, the star product on Kahler manifolds, and Topological Quantum Field Theory
The formula for the star product of two functions and on the space of connections arising from a functional integral in Chern-Simons gauge theory where translates of and are ”stacked on top of each other” is not unexpected from the point of view of geometric quantization and topological quantum field theory.
To see this, we first recall the construction of the star product for Kahler manifolds [4, 3, 11, 17]. Let be a compact Kahler manifold, and let be a Hermitian holomorphic line bundle of curvature The geometric quantization of (where ) is given in this Kahler polarization by the Hilbert space
of holomorphic sections of this space is a finite dimensional subspace of the infinite dimensional Hilbert space of sections of 222222The metric arises from the Hermitian metric on and the symplectic volume form on Let denote the projection.
Given , we obtain a Toeplitz operator on by defining
for any These Toeplitz operators generate a non-commutative, associative algebra, and we have [4, 3, 11, 17] for asymptotically as
where is a star product on
Now let be a compact Riemann surface and let be a compact Lie group; take for simplicity We may consider the space as a space of stable bundles on the Riemann surface This is a Kahler variety in its own right, and possesses a Hermitian line bundle of curvature given by the Kahler form.232323Depending on the precise details, the variety may be singular; we do not expect the singularities to affect any of these considerations. It may therefore be quantized as above, and we may apply (5.1) to obtain a star product on the algebra of smooth functions on Among these functions are the Wilson loops.
On the other hand, the quantization is part of a three dimensional topological quantum field theory, giving the Hilbert space associated by the theory to the two manifold underlying the Riemann surface One may reasonably hope that this topological quantum field theory would extend to three-manifolds with immersed Wilson loops. Given a smooth function we would then expect to obtain an element of –that is, an operator –by placing the Wilson loops composing in a ”block” or cobordism One may also reasonably hope, from the axioms of Topological Quantum Field Theory, that multiplication of two such operators would amount to composing the corresponding cobordisms; that is, to ”stacking the blocks”.242424There is a vast literature on Skein Modules related to this type of construction in Topological Quantum Field Theory. (See Figure 1.) We would expect these operators to coincide with the Toeplitz operators in the Kahler quantization The product obtained by ”stacking the blocks” is then the operator product, and hence, by (5.1), gives the star product. Thus ”stacking” the two functions and to obtain the star product of and as we did in Definition 4.4, is precisely what would be predicted by Topological Quantum Field Theory, Geometric Quantization, and the relation between them in the case of moduli spaces.
In our Chern Simons functional integral in axial gauge, something like this rather subtle picture252525Subject to issues of gauge invariance and the Second Reidemester move; see Section 5.2. is obtained from the simple relation between and We hope this serves as an illustration of the power of functional integrals when they can be treated mathematically.
5.2. Gauge invariance, quantization of moduli spaces, and the Second Reidemeister move.
We saw in Section 2 that our Poisson bracket on the Wilson loops on the space of connections gave rise to Goldman’s Poisson bracket on these same Wilson loops, considered as functions on the moduli spaces of flat connections on these moduli spaces are the character varieties considered above. This follows from the general fact that on a Hamiltonian space the Poisson bracket of invariant functions gives a Poisson bracket on the algebra of functions on the symplectic quotient coinciding with the Poisson bracket coming from the symplectic form on the quotient.
Fedosov [8] shows that the deformation quantization of a Hamiltonian -space also descends to the quotient in a similar way. We may consider what this would mean for the case of the space of connections and its quotient by the gauge group. Given two homotopic loops on there is no necessary relation between the corresponding Wilson loops, considered as functions on But as functions on the quotient two homotopic loops give rise to identical functions. So if our star product of such functions on were to descend to a star product on the quotient, the star product of loops would have to be a homotopy invariant. In particular, it would have to be invariant under the slide move in Figure 2—the two dimensional analog of the Second Reidemeister move for knot diagrams.
Now in [18] we showed that in the case of the expectation was not invariant under the Second Reidemeister move; extracting a link invariant involved a shift in the value of the expectation of a single unknotted loop from the value expected in the Chern-Simons path integral, which is 262626In [18] we used a slightly different normalization for the expectation of Wilson loops associated with unoriented curves when which would change this expectation by a sign. Since this value is also the value of the Wilson loop corresponding to a nullhomotopic simple closed curve, considered as a function on the space of flat connections, the same issue will affect the star product as well. So our star product on will not give a star product on but the difficulty is due to a shift in something like a ground state energy.
From the point of view of [8], there is a geometric explanation of why we do not obtain a star product on the quotient. Fedosov’s construction of the deformation quantization of a symplectic manifold thought of as a star product on functions, involves a choice of a symplectic connection on While the resulting algebra, thought of in terms of a sheaf of algebras on is in an appropriate sense independent of this choice, this invariance is not so straightforward to describe in terms of the star product on functions.
Now in our construction of the star product on we made a particular choice of symplectic connection, arising from the affine structure of This connection is not gauge invariant, and hence cannot be expected to yield a star product on
Turning our attention back to Chern-Simons gauge theory, and motivated by the observation that gauge (non-)invariance may be behind the failure of the star product to be invariant under homotopy, it may be interesting to consider whether a more sophisticated treatment of the time independent gauge transformations in Chern Simons gauge theory in axial gauge may modify expectations in the theory sufficiently so as to obtain invariance under the Second Reidemeister move, and give a geometric explanation for the shifted expectation

### v2

@@
 In our case this relation arises from explicit algebraic formulas arising from the (mathematically well-defined) functional integrals of [33].
-***PRELIMINARY***
 Contents
+- 1 Introduction
+- 2 The Poisson Bracket
+- 3 Chern Simons gauge theory in axial gauge on
+- 4 Chern Simons gauge theory in three dimensions and the star product on the space of connections in two dimensions
+- 5 Some Comments and Speculations: Relation to Geometric Quantization; Quantization of Moduli spaces
+- A Some comments and speculations on a coproduct structure
 1. Introduction
@@
 for where is the Poisson bracket on
-The construction of the star product for symplectic manifolds in general was given by deWilde and Lecomte [6] and by Fedosov [7], and for general Poisson manifolds by Kontsevich [15].
+The construction of the star product for symplectic manifolds in general was given by deWilde and Lecomte [10] and by Omori, Maeda, and Yoshioka [28] and Fedosov [12], and for general Poisson manifolds by Kontsevich [23].
 The resulting star products are often not so easy to compute.
@@
 This parallels the relation we found in this paper between deformation quantization of the space of connections and Chern-Simons gauge theory.
-We close by discussing the relation between the (non)-invariance of the expectations in the Chern-Simons gauge theory of [18] under the Second Reidemeister move and the (non)-invariance of the star product on under reduction.
+We next discuss the relation between the (non)-invariance of the expectations in the Chern-Simons gauge theory of [33] under the Second Reidemeister move and the (non)-invariance of the star product on under reduction.
 While the lack of invariance of expectations in Chern-Simons gauge theory is mysterious, the non-invariance of the star product arises from a known cause—the gauge non-invariance of the affine symplectic connection on So this hints at a possible resolution of both issues, as mentioned above.
+Finally we remark briefly on the known relation between skein modules and quantized Teichmuller space.
+From our point of view this provides evidence that the issues discussed above are solvable.
+Our results would then provide a geometric proof, based on mathematical functional integrals, of the relation between Chern Simons gauge theory and deformation quantization of moduli spaces (such as Teichmuller space).
+We also ask whether the coincidence we see between the results for (related to quantum invariants or the Kauffman bracket) and those for (related to Teichmuller space) would give some insight into the volume conjecture.
+We close with an appendix where we construct various coproducts on the functions we have studied, and ask whether any of those are related to Turaev’s coproduct.
 2. The Poisson Bracket
@@
 Turning our attention back to Chern-Simons gauge theory, and motivated by the observation that gauge (non-)invariance may be behind the failure of the star product to be invariant under homotopy, it may be interesting to consider whether a more sophisticated treatment of the time independent gauge transformations in Chern Simons gauge theory in axial gauge may modify expectations in the theory sufficiently so as to obtain invariance under the Second Reidemeister move, and give a geometric explanation for the shifted expectation
+5.3. Skein modules and quantized Teichmuller space; possible relation to the volume conjecture
+A relation similar to the one we have found between quantization of the space of connections and the Kauffman bracket is the relation between skein modules and the deformation quantization of Teichmuller spaces of [14, 21].
+A few references on this relation are [32, 4, 7, 8, 24, 29].
+Since Teichmuller space occurs as a component of the reduction of the space of connections by the gauge group, where and since as remarked in [33] and above, the expectations in Chern Simons gauge theory are the same for and our methods should give a geometric proof of this relation using mathematical functional integrals, if we are able to find a gauge invariant symplectic connection, or at least a modification of the Chern Simons functional integral which is invariant under the second Reidemeister move, as explained in Section 5.2.
+We should note that results on the uniqueness of star products give reason to expect that the desired star product, as a star product on the space of connections, is equivalent to the star product in this paper.
+Recall that given a star product on a symplectic manifold and a differential operator on of the form
+where is another differential operator on a new star product can be defined by
+for two such star products are called equivalent.
+Then [3, 11, 18] if any two star products are related by a combination of equivalence and a change of the parameter 272727See [18], Corollary 9.5.
+Since is contractible (though infinite dimensional), we may hope for a similar result here.
+In a yet more speculative vein, the relation between the Kauffman bracket and quantized Teichmuller space, and more generally the coincidence of the expectations of loops in Chern Simons gauge theory for the group with those for Chern-Simons gauge theory for the groups and may be related to the volume conjecture [20, 25, 26]
+Appendix A Some comments and speculations on a coproduct structure
+In [2] Turaev showed that the Goldman bracket on the algebra of oriented loops on a two manifold could be supplemented with a cobracket, forming a bialgebra structure.
+Something like this may be visible in our gauge theoretic construction.
+In particular, if we consider the case and take , the monomials in the Wilson loops might be expected to be linearly independent, and form a basis for a vector space of functions on If we take the duals of those functions to be a basis for the dual vector space then we obtain a formula similar to (2.12), giving a bracket This bracket is equivalent to a cobracket on which is the Turaev cobracket.
+The bracket on can be exponentiated to give an associative product on in line with the formulas in Definition 4.4 and given more explicitly in Section 4.1. This in turn gives a coproduct on
+Question A.1.
+Do the product and coproduct combine to make the loops on a compact oriented two manifold into a Hopf Algebra??
+If we could overcome the difficulties of gauge invariance, we might expect to obtain also a product and coproduct on the homotopy classes of loops on a compact oriented Riemann surface; equivalently, on the Wilson loops, considered as functions on In this case, there is a unit (consisting of the trivial loop) and counit (consisting of its dual), and likewise an inverse (by reversing the orientation of the loop) and antipode.
+We may ask whether these fit into any kind of Hopf algebra structure.
+Question A.2.
+Is there an analogous construction for unoriented loops?
+Of course all of these speculations ignore transversality issues and other troubles, which mean that none of our products or coproducts are really well-defined.
+They are very far from established facts.
+A.1. A geometric coproduct
+There is also another coproduct construction on the algebra of loops, arising from Bott periodicity. Again, let us take
+The space is a contractible space on which the group of based maps from to acts freely.
+The functions are -invariant functions on this space.
+We can therefore consider them as functions on the classifying space Note that if we are given a curve in and a point of we obtain by composition an element of taking the trace we get a complex number, and the resulting function on (or, equivalently, -invariant function on is the Wilson loop If we take these are functions on where is the space of based loops on which are homotopically trivial, by Bott periodicity.
+The space is a topological group (from the group structure on ) as well as an -space (from concatenation of loops).
+This means that the space also has two space structures.
+So the space of functions on acquires two coproduct structures, one of which is coassociative and one of which is coassociative ”up to homotopy”.
+A third construction of a coproduct arises from a map we can write directly: Identifying with the Grassmannian, consider the map arising from the map taking two -dimensional subspaces of Hilbert space to Taking the limit we obtain a map And this again gives a type of product structure on again giving a coproduct on the functions on 282828This construction on is well known (see e.g.
+[1]), but its application to is not, to my knowledge.
+Note that all these coproducts are coassociative, but not co-commutative; and they do not come in a parametrized family in any obvious way.292929It is possible to introduce an integer parameter into the third construction by modifying the map to
+Question A.3.
+Do any of these geometric coproducts have anything to do with the Turaev cobracket?